LHC 2012 - latest news

Week 38: Machine coordinators: Gianluigi Arduini, Mike Lamont

Plan for the week: Technical stop and recovery

Date

Summary

Link to report from morning meeting

Monday 24 September
  • Cryo back at 1:00
  • Powering tests R78/R81
  • 05:20 Beam in
  • 08:00 Pilot ramp & squeeze

8:30 meeting
(W38 summary)

Cryogenics report
(Serge Claudet)

Sunday 23 September
  • Cryo recovery
09:00 meeting
Saturday 22 September
  • Cryo recovery (pt. 2 and pt. 8)
09:00 meeting

Friday 21 September

  • Technical stop
  • 11:24 electrical cut at Pt 2 (before the re-start of the compensator) caused loss of cryogenics at point 2 including cold compressor
  • 12:00 Compensator in point 2 restarted
  • 14:00 MKI 8 ready for operation.
  • 16:00 trip of cryogenics in point 8: 24 V power supply causing an emergency stop of the cryogenics
  • 17:00 SIG informed that the intervention on the electrical network is postponed to next week
  • 18:20 Machine patrolled. Pressures in the BSRTs in the 10-9 10-8 mbar range. RF available.
  • 18:00 20:30 Beams to last TED for Q20 optics extraction verification successful
  • 20:20 LHC closed after intervention on 60 A power converter
  • 23:00 Problems with heaters controls of the cryogenics cold box in point 8.
  • Cryo recovery  (pt. 2 and pt. 8)
08:30 meeting

Thursday 20 September

  • Technical stop
 
Wednesday 19 September
  • Technical stop
 
Tuesday 18 September
  • Technical stop
 
Mon 17 September
  • 06:03 Dumped fill #3071, integrated lumi 143 pb-1
  • Start of TS3
08:30 meeting